OpenStudy (anonymous):

Find a power series representative of the function and determine the interval of convergence. \[f(x)=\frac{3}{1-x^4}\]

OpenStudy (anonymous):

sorry it's the first problem in the section, can you give me a hint on where to start?

OpenStudy (anonymous):

well I understand that I have to change it into the form \[\sum\] somehow

OpenStudy (anonymous):

\[\frac{1}{1-x^4}=\sum_{k=0}^{\infty}x^{4k}\] how and why does the left side equal the right side?

OpenStudy (anonymous):

I have access to the solutions, but don't quite understand them

OpenStudy (anonymous):

I can figure out how to manipulate \[\frac{1}{1-x}= \sum_{n=0}^{\infty} x^n\] to look like \[\frac{1}{1-x^4}=\sum_{k=0}^{\infty}x^{4k}\] but I would prefer to understand how so

OpenStudy (anonymous):